OpenStudy (anonymous):

simplify the square root of (x + 8) squared

OpenStudy (ranga):

Is the expression \[\Large \sqrt{(x+8)^2}\]?

OpenStudy (ranga):

If it is, then the square root and the square will cancel each other out. \[\Large \sqrt{A^2} = (A^{2})^{1/2} = A^{2*1/2} = A^1 = A\]